For what value(s) of k, if any, will the system have no solution, a unique solution, and infinitely many solutions? (Enter your answers as a comma-separated list. If an answer does not exist, enter DNE.)


x − 2y + 8z = 6

x + y + z = k

2x − y + 9z = k2

1
Expert's answer
2021-04-19T07:26:37-0400

Consider the system of linear equations

x-2y+8z=6

x +y +z=k

2x-y+9z=k2


The augmented matrix for the above linear system:


A = "\begin{bmatrix}
 1 & -2 & 8 | & 6 \\
 1 & 1 & 1 | & k \\
 2 & -1& 9 | & k^2
\end{bmatrix}"


The row reduction process as follows:


"R_2\implies R_2-R_1"


A = "\begin{bmatrix}
 1 & -2 & 8&| & 6 \\
 0 & 3 & -7&| & k-6 \\
 2 & -1& 9&| & k^2
\end{bmatrix}"


"R_3\implies R_3-2R_1"


A = "\begin{bmatrix}
 1 & -2 & 8&| & 6 \\
 0 & 3 & -7&| & k-6 \\
 0 & 3& -7&| & k^2-12
\end{bmatrix}"


"R_3\implies R_3-R_2"


A = "\begin{bmatrix}
 1 & -2 & 8&| & 6 \\
 0 & 3 & -7&| & k-6 \\
 0 & 0& 0&| & k^2-k-6
\end{bmatrix}" ........ (1)


The system has no solution if the the rank of the matrix "\begin{bmatrix}
 1 & -2 & 8 \\
 0 & 3 & -7 \\
 0 & 0& 0
\end{bmatrix}" is not equal to rank of matrix "\begin{bmatrix}
 1 & -2 & 8&| & 6 \\
 0 & 3 & -7&| & k-6 \\
 0 & 0& 0&| & k^2-k-6
\end{bmatrix}"


Let "k^2-k-6\ne 0"

"(k+2)(k-3)\ne0"

"k\ne-2,3"

Thus, the rank of matrix "\begin{bmatrix}
 1 & -2 & 8 \\
 0 & 3 & -7 \\
 0 & 0& 0
\end{bmatrix}" is 2.


Thus, the rank of matrix "\begin{bmatrix}
 1 & -2 & 8&| & 6 \\
 0 & 3 & -7&| & k-6 \\
 0 & 0& 0&| & k^2-k-6
\end{bmatrix}" is 3.


Hence the system has no solution if "k\ne-2,3"

The system has unique solution if the rank of matrix "\begin{bmatrix}
 1 & -2 & 8 \\
 0 & 3 & -7 \\
 0 & 0& 0
\end{bmatrix}" is equal to rank of matrix "\begin{bmatrix}
 1 & -2 & 8&| & 6 \\
 0 & 3 & -7&| & k-6 \\
 0 & 0& 0&| & k^2-k-6
\end{bmatrix}" "=3"


This is not possible.

So, the system has no unique solution.


The system has infinitely many solutions if the rank of matrix "\begin{bmatrix}
 1 & -2 & 8 \\
 0 & 3 & -7 \\
 0 & 0& 0
\end{bmatrix}" is equal to rank of matrix "\begin{bmatrix}
 1 & -2 & 8&| & 6 \\
 0 & 3 & -7&| & k-6 \\
 0 & 0& 0&| & k^2-k-6
\end{bmatrix}" less than 3.


Let "k^2-k-6=0"

"(k+2)(k-3)=0"

"k=-2,3"

Thus, the rank of matrix "\begin{bmatrix}
 1 & -2 & 8 \\
 0 & 3 & -7 \\
 0 & 0& 0
\end{bmatrix}" is 2.



Thus, the rank of matrix "\begin{bmatrix}
 1 & -2 & 8&| & 6 \\
 0 & 3 & -7&| & k-6 \\
 0 & 0& 0&| & k^2-k-6
\end{bmatrix}" is 2.


Hence, the system has infinitely many solutions if "k=-2,3" .